OpenStudy (anonymous):

Let j = 16. What is the value of k in the equation k = 2 × j + 4?

Mehek (mehek14):

\(k = 2 * 16 + 4\) what's 2 * 16

OpenStudy (anonymous):

36

Mehek (mehek14):

@ian_s giving direct answers is not allowed read the coc if you want to know more

OpenStudy (anonymous):

you guys are ALOT smarter than me...

Mehek (mehek14):

use PEMDAS

Mehek (mehek14):

there are no parenthesis nor exponents but there is multiplication

OpenStudy (anonymous):

We're also a lot older, so we've had a lot more time to work with this stuff. So no need to worry, it takes time to get it for everyone :)

Mehek (mehek14):

so multiplication comes first

Mehek (mehek14):

\(2*16 = \)

OpenStudy (anonymous):

legendarysadist love the name and @Mehek14 thanks you guys are awsome

Mehek (mehek14):

did you get the answer?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

is it 32?

OpenStudy (anonymous):

bye @LegendarySadist

OpenStudy (anonymous):

2*16= 32 then you add 4.